OpenStudy (anonymous):

Find the derivative of 8 / (2+3e^(-x/2))

OpenStudy (anonymous):

move the denominator to numerator, and use a power rule, or quotation rule

OpenStudy (anonymous):

Could you show how to move the denom. to the num.?

OpenStudy (anonymous):

Use the quotient rule. \[\frac{d}{{dx}}\left( {\frac{{f\left( x \right)}}{{g\left( x \right)}}} \right) = \frac{{\frac{d}{{dx}}f\left( x \right)g\left( x \right) - f\left( x \right)\frac{d}{{dx}}g\left( x \right)}}{{g^2 \left( x \right)}}\] Remember when differentiating the exponential term to use the chain rule!

OpenStudy (anonymous):

Moving the denominator to the numerator is not a good idea since the denominator is a linear combination. Moving the denominator to the numerator only makes sense when the denominator is a product.

OpenStudy (anonymous):

misplaced comma

OpenStudy (anonymous):

understand that 1/x is the same as x^-1 by this knowledge, you can make use of the product rule

OpenStudy (anonymous):

Could you show me what you got for the derivative please?

OpenStudy (anonymous):

well well... could you show me your attempt at solving this problem?

OpenStudy (anonymous):

i have no idea of what to do

OpenStudy (anonymous):

my my.... then you have to go back from the very beginning

OpenStudy (anonymous):

how about you read about the tangent problem, limits and then we will discuss this further

OpenStudy (anonymous):

The denominator CANNOT be moved the the numerator as the denominator is a linear combination of terms.

OpenStudy (anonymous):

that is some pretty BIG suggestion you have, young man

OpenStudy (anonymous):

Do not move the denominator to the numerator. Use the quotient rule which I typed above. It is not possible to move a linear combination like that to the numerator. It MUST be a product of terms, not a sum.